the squares of 2 consecutive odd numbers differ by 88, what are the numbers?
Answer by RAY100(1637)

Let x,,,,x+2 ,,,,be consecutive odd numbers
.
(x+2)^2 -(x)^2 =88
.
x^2 +4x +4 -x^2 =88
.
4x =84
.
x=21
.
Numbers are 21,,23
.
.
check
23^2-21^2 = 88,,,,,,ok
.